OpenStudy (anonymous):

Solve the posted equation. Show work please. Thank you! (:

OpenStudy (anonymous):

\[\sqrt{3x+5}=\sqrt{4x-2}\]

OpenStudy (anonymous):

square both sides, get \[3x+5=4x-2\] solve for x

OpenStudy (chriss):

So if you square both sides, you keep both sides equal, but you no longer have to deal with the square roots so you get the simpler problem, 3x+5=4x-2

OpenStudy (anonymous):

The answer is 7 but my answer was x+7 Thats why i'm confused. :/

OpenStudy (anonymous):

my answer was -x+7

OpenStudy (chriss):

Well this problem is completely solvable for x, so you should be able to completely isolate x to one side. 3x+5=4x-2 I would begin by adding 2 to both sides.

OpenStudy (anonymous):

\[3x+5=4x-2\] \[5=x-2\] \[7=x\]

OpenStudy (anonymous):

:O Thank you smart people!

OpenStudy (chriss):

your welcome